If the price of a commodity increases from $.40 and $. 60 and its supply increases from 6 kg and 9 kg . Calculate its elasticity of supply.
1
Expert's answer
2020-09-28T10:05:13-0400

Elasticity of Supply ="\frac{\% Change in Quantity} {\% Change in price}"


Percentage change in Price =40"-" 60=20

="\frac{20}{60}×100 =33.33\%"


Percentage change in Quantity =6"-" 9=3

="\frac{3}{6} ×100 =50\%"


Thus ;Elasticity of Supply ="\frac{50\%}{33.33\%} =1.5"


Elasticity of Supply =1.5